OpenStudy (anonymous):

Prove that sinθ.sin(60 - θ).sin(60 + θ) = 1/4 sin 3θ .

OpenStudy (anonymous):

use d identity 2sinAsinB=cos(A-B)-cos(A+B)

OpenStudy (anonymous):

use it where ?

OpenStudy (anonymous):

or use d identity sin(a+b) *Sin(a-b)= sin^2a- Sin^2b

OpenStudy (anonymous):

used the 1st identity for last two terms i.e. Sin(60+x)*Sin(60-x)

OpenStudy (anonymous):

1st multiply d num n denom by 2

OpenStudy (anonymous):

is there a dr ?

OpenStudy (anonymous):

then 2Sin(60+x)*Sin(60-x)=cos2x-cos120

OpenStudy (anonymous):

then multiply it by sinx

OpenStudy (anonymous):

u'll have (sinxcos2x-sinxcos120)/2

OpenStudy (anonymous):

den cos2x=1-2sin^2x and simplify

OpenStudy (anonymous):

and then .. i don't know akash after 

OpenStudy (anonymous):

and is that correct ?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

and then what to do ?

OpenStudy (anonymous):

use co120=-1/2 and sin3x=3sinx-4sin^3x

OpenStudy (anonymous):

cos120=-1/2...just u need to simplify and arrange the terms

OpenStudy (anonymous):

try it....

OpenStudy (anonymous):

yea .. okay got it , thanks @akash123 :)

OpenStudy (anonymous):

good...:)